# Isotropic cosmic birefringence from string axion domain walls without cosmic strings, and DESI results

### v2

@@ 1 Introduction
where $c_{\gamma}$ is an anomaly coefficient, [^1] Note that the definition of $c_{\gamma}$ differs from the definition of the electromagnetic anomaly coefficient $E$ for the QCD axion adopted in some literature by a factor of $2$. Our definition of $c_{\gamma}$ follows the nomalization of the U(1) anomaly used in, e.g., Ref. [10]. $\alpha$ is the fine structure constant, and $F_{\mu\nu}$ and $\tilde{F}_{\mu\nu}$ denote the field strength and its dual, respectively. In the literature, an axion with such a photon coupling is often referred to as an axion-like particle (ALP). For comprehensive reviews on axions and related topics, see Refs. [11, 12, 13, 14, 15, 16, 10]. The axion photon coupling is constrained as $g_{\phi\gamma\gamma}<6.3\times 10^{-13}\,\mathrm{GeV}^{-1}$ or equivalently $f_{\phi}>3.7\times 10^{9}c_{\gamma}$ GeV for axions lighter than $10^{-12}$ eV [17].
@@
If the axion initially resides near the hilltop of the potential, it begins to evolve when the Hubble parameter becomes comparable to its mass. [^2] The anharmonic effect, which modifies the axion dynamics near the hilltop, will be discussed later. At this stage, two possibilities arise. One possibility is that fluctuations of the axion do not overcome the potential maximum, so the field remains nearly homogeneous and begins to oscillate around one of the potential minima. In this case, the isotropic CB is predicted by Eq. (9). The other possibility is that axion fluctuations cross the potential maximum, leading to the population of both vacua. As a result, domain walls form between these vacua.
In particular, it was proposed by [^3] Two of the present authors (FT and WY). Ref. [30] that the reported value of the rotation angle can be explained by ALP domain walls in the scaling solution. In this scenario, the region around Earth lies within a single domain, while the domain wall network at recombination contains $10^{3\text{--}4}$ domains across the LSS. As a result, the axion field value on the LSS effectively averages to the midpoint between two adjacent vacua. Depending on which vacuum the axion occupies in a given domain, the rotation angle $\Delta\Phi(\Omega)$ takes one of two possible values—either 00 or $0.42\,c_{\gamma}$ degrees. Consequently, each domain contributes one bit of information to the cosmic birefringence, and this phenomenon has been termed kilobyte cosmic birefringence (KBCB). Remarkably, spatial averaging over many domains yields an isotropic CB prediction that is nearly identical to that in the case where the axion starts uniformly near the hilltop and evolves coherently, as given by Eq. (9).
@@
On the other hand, the Dark Energy Spectroscopic Instrument (DESI) DR2 baryonic acoustic oscillation measurements [61, 62], based on a larger and higher-quality dataset than DR1 [63], improve the statistical precision by about $40\%$. Combined with CMB and supernova data [64, 65], these results provide stronger hints—at the $2.8\sigma$ to $4.2\sigma$ level—for a time-dependent dark energy component, beyond a simple cosmological constant.
Various explanations for this possibility have been proposed [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. Among them, models of time-varying dark energy motivated by cosmic birefringence have been studied in Refs. [66, 67, 75] (see also Ref. [33] before the DESI data release and earlier studies [79, 80]). Such time-dependent dark energy may be connected to the cosmological constant problem [81, 67], and indeed, it arises naturally in models that address it [81]. In this context, a hilltop axion model with matter effects to also address the $S_{8}$ tension has been proposed [82] (see also [40] for domain wall formation and [83, 84, 85, 86, 87, 88] for general hilltop axion models).
@@ 2 String axion domain walls and isotropic cosmic birefringence
In this section, we investigate a scenario in which an axion initially located near the hilltop of its potential leads to the formation of domain walls without strings, and explore the resulting implications for cosmic birefringence. For our purposes, it is sufficient to consider the following potential:
@@
where $\Lambda$ is a dynamical scale, $m_{\phi}\equiv\Lambda^{2}/f_{\phi}$ is the mass of the axion. The second line represents an expansion around the origin. This potential is invariant under the discrete shift symmetry transformation (1), and consequently admits degenerate vacua. One of the maxima is located at the origin, and there are degenerate vacua at $\phi=\pm\pi f_{\phi},\pm 3\pi f_{\phi},\cdots$. In the following we consider a case in which $\phi$ is initially around the origin so that the dynamical range of $\phi$ is limited to $-2\pi f_{\phi}<\phi<2\pi f_{\phi}$.
@@
Note that the anharmonic effect could delay the onset of oscillations.
Suppose that $\phi_{0}$ is sufficiently close to the hilltop at the origin such that
@@
Figure 1: Redshift of domain wall formation estimated using Eq. (11) for $\sigma_{\rm DW}=8m_{\phi}f_{\phi}^{2}=(0.1\,\mathrm{MeV})^{3}$, $(1\,\mathrm{MeV})^{3}$, and $(10\,\mathrm{MeV})^{3}$ from left to right. See Fig. 2 for the delay due to the anharmonic effect.
[…]
Domain walls form shortly after the axion begins to oscillate, at the time when condition (11) is satisfied. The corresponding redshift at the onset of oscillations, estimated using this condition, is shown in Fig. 1. Here we use the fact that the domain wall tension is given by $\sigma_{\rm DW}\simeq 8m_{\phi}f_{\phi}^{2}$ for the potential (10) and fix the tension by $\sigma_{\rm DW}=(0.1\,\mathrm{MeV})^{3}$, $(1\,\mathrm{MeV})^{3}$, and $(10\,\mathrm{MeV})^{3}$, from left to right. Note that this estimate is valid when the field fluctuations are not much smaller than $f_{\phi}$ or when the zero mode is not very close to the hilltop. If the fluctuations are small and the zero mode is in the vicinity of the hilltop, the anharmonic effect becomes significant and delays the onset of oscillations. Figure 2 illustrates how the anharmonic effect causes this delay. In this sense, the redshift shown in Fig. 1 should be interpreted as the maximal redshift of the domain wall formation.
@@
where $\bar{\phi}_{\rm LSS}$ denotes the angular average of $\phi_{\rm LSS}(\Omega)$ over the LSS, and we have used the fact that $N_{e}\simeq 50\sim 60$. Since the observer resides in one of the two vacua after domain wall formation, the isotropic cosmic birefringence is given by
@@
where the positive sign in the second expression is chosen to match the observed sign of the rotation angle.
Note that the origin of the isotropic CB in this scenario differs from that in the original KBCB mechanism. There, domain wall formation has already occurred by the time of last scattering, and both adjacent vacua are populated on the LSS. The average axion field is given by $\bar{\phi}_{\rm LSS}=(\phi_{\rm min,1}+\phi_{\rm min,2})/2$, where $\phi_{\rm min,1}$ and $\phi_{\rm min,2}$ denote the field values of the adjacent minima, one of which corresponds to our local vacuum. The isotropic CB then arises from averaging over these domains, while anisotropic CB is also generated due to spatial variations in the field. In this case, the predicted rotation angle may carry uncertainty due to the population bias. In contrast, in the present scenario, the axion remains near the hilltop until well after recombination, and the field on the LSS is nearly homogeneous. Consequently, the CB from CMB photons is purely isotropic and free from ambiguity related to population bias. Remarkably, despite the differences in the underlying dynamics, the predicted isotropic rotation angle is the same in both scenarios.
@@
We also note that in this scenario, the formation time of domain walls can be significantly later than the maximal estimate shown in Fig. 1, if the fluctuations are sufficiently small. This behavior is illustrated in Fig. 2, where we plot the evolution of a homogeneous axion field $\phi$, representing a typical displacement from the potential maximum. Here, we numerically solve the equation of motion
@@
[…]
We see that, in order to avoid the domain wall problem for a string axion, the domain walls must form after the recombination era. As a result, we do not expect anisotropic birefringence in the CMB. However, if the walls form well before reionization, photons from that epoch can still experience anisotropic birefringence. Current observational constraints on such anisotropic cosmic birefringence remain consistent even with scenarios involving cosmic strings [95].
In addition, if the domain walls formed significantly earlier than the present time, they could have emitted gravitational waves [96]. Such gravitational waves are searched for alongside CMB anisotropies induced by the gravitational potential of domain walls [89, 97, 98, 99].
If the domain walls have only just formed today, they would not have reached the scaling regime. In that case, many closed domain walls could exist within the current horizon. This is possible if the wall thickness is much smaller than the horizon size, which can occur when the onset of oscillations is significantly delayed by the anharmonic effect near the hilltop. Photons from galaxies at relatively low redshift may experience a polarization rotation if they cross domain walls an odd number of times. The resulting rotation angle is predicted to be $0.42\,c_{\gamma},\mathrm{deg}$, which is twice the isotropic birefringence angle observed in the CMB.
@@ 3 Time-Varying Dark Energy
If domain walls are formed, however, it becomes challenging to simultaneously satisfy the constraint on their tension—imposed by the requirement that their energy density remain below roughly $10^{-5}$ of the present total energy density—and realize a significant deviation of the dark energy equation of state from that of a cosmological constant ($w=-1$). Although it is in principle possible to suppress the energy stored in the domain walls while enhancing the energy in the bulk oscillations of the axion field, which we will comment on later, we here focus on the implications of a hilltop initial condition for a string axion—suggested by the isotropic CB—for the DESI DR2 observations, assuming that domain walls are not formed.
@@
In Fig.3, we show the redshift dependence of the dark energy equation of state parameter $w_{\rm DE}(z)$ in the hilltop axion scenario under consideration. For the red (blue) line, we set $f_{\phi}=10^{16}\,\mathrm{GeV}$, $\phi_{\rm ini}=3\,\mathrm{GeV}$, and $m_{\phi}=1.13\times 10^{-31}\,\mathrm{eV}$ ($f_{\phi}=6\times 10^{16}\,\mathrm{GeV}$, $\phi_{\rm ini}=10^{15}\,\mathrm{GeV}$, and $m_{\phi}=1.89\times 10^{-32}\,\mathrm{eV}$), and compare it with the data point from DESI DR2[61]. Our prediction remains consistent with the $\Lambda$CDM model for $z>1$, showing no significant deviation in that range. [^5] To obtain a better fit at $z>1$, modifications to the matter or radiation sector may be required, which are model-dependent. We also note that the $z<1$ bin has the smallest error bar [61].
@@
Figure 3: Redshift dependence of the dark energy equation of state parameter $w_{\rm DE}(z)$ in the hilltop axion scenario. The red (blue) line corresponds to $f_{\phi}=10^{16}\,\mathrm{GeV}$, $\phi_{\rm ini}=3\,\mathrm{GeV}$, and $m_{\phi}=1.13\times 10^{-31}\,\mathrm{eV}$ ($f_{\phi}=6\times 10^{16}\,\mathrm{GeV}$, $\phi_{\rm ini}=10^{15}\,\mathrm{GeV}$, and $m_{\phi}=1.89\times 10^{-32}\,\mathrm{eV}$). The black point shows the DESI DR2 result from Ref. [61].
@@
where we define $t_{\mathrm{osc}}$ by $t_{\mathrm{osc}}\equiv 2/(3H_{\Lambda\mathrm{CDM}}(z_{\mathrm{osc}}))$ with $H_{\Lambda\mathrm{CDM}}(z_{\mathrm{osc}})\equiv H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z_{\mathrm{osc}})^{3}}$. From these conditions, we obtain the required tuning of the initial condition for fixed values of $f_{\phi}$ (or $m_{\phi}$) as shown in Fig. 4. We also show the parameters from Fig. 3 by red and blue dots in Fig. 4, which roughly match the analytical estimate.

### v3

@@ 1 Introduction
where $c_{\gamma}$ is an anomaly coefficient, [^1] Note that the definition of $c_{\gamma}$ differs from the definition of the electromagnetic anomaly coefficient $E$ for the QCD axion adopted in some literature by a factor of $2$. Our definition of $c_{\gamma}$ follows the normalization of the U(1) anomaly used in, e.g., Ref. [10]. $\alpha$ is the fine structure constant, and $F_{\mu\nu}$ and $\tilde{F}_{\mu\nu}$ denote the field strength and its dual, respectively. In the literature, an axion with such a photon coupling is often referred to as an axion-like particle (ALP). For comprehensive reviews on axions and related topics, see Refs. [11, 12, 13, 14, 15, 16, 10]. The axion photon coupling is constrained as $g_{\phi\gamma\gamma}<6.3\times 10^{-13}\,\mathrm{GeV}^{-1}$ or equivalently $f_{\phi}>3.7\times 10^{9}c_{\gamma}$ GeV for axions lighter than $10^{-12}$ eV [17].
@@
If the axion initially resides near the hilltop of the potential, it begins to evolve after the Hubble parameter becomes comparable to its mass. [^2] The anharmonic effect, which modifies the axion dynamics near the hilltop, will be discussed later. At this stage, two possibilities arise. One possibility is that fluctuations of the axion do not overcome the potential maximum, so the field remains nearly homogeneous and begins to oscillate around one of the potential minima. In this case, the isotropic CB is predicted by Eq. (9). The other possibility is that axion fluctuations cross the potential maximum, leading to the population of both vacua. As a result, domain walls form between these vacua.
In particular, it was proposed by [^3] Two of the present authors (FT and WY). Ref. [30] that the reported value of the rotation angle can be explained by ALP domain walls in the scaling solution. In this scenario, the region around Earth lies within a single domain, while the domain wall network at recombination contains $10^{3\text{--}4}$ domains across the LSS. As a result, the axion field value on the LSS effectively averages to the midpoint between two adjacent vacua. Depending on which vacuum the axion occupies in a given domain, the rotation angle $\Delta\Phi(\Omega)$ takes one of two possible values—either 0 or $0.42\,c_{\gamma}$ degrees. Consequently, each domain contributes one bit of information to the CB, and this phenomenon has been termed kilobyte cosmic birefringence (KBCB). Remarkably, spatial averaging over many domains yields an isotropic CB prediction that is nearly identical to that in the case where the axion starts uniformly near the hilltop and evolves coherently, as given by Eq. (9).
@@
On the other hand, the Dark Energy Spectroscopic Instrument (DESI) DR2 baryonic acoustic oscillation measurements [61, 62], based on a larger and higher-quality dataset than DR1 [63], improve the statistical precision by about $40\%$ as stated in Sec. VI of Ref. [61]. Combined with CMB and supernova data [64, 65], these results provide stronger hints—at the $2.8\sigma$ to $4.2\sigma$ level—for a time-dependent dark energy component, beyond a simple cosmological constant.
Various explanations for this possibility have been proposed [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. Among them, models of time-varying dark energy motivated by CB have been studied in Refs. [66, 67, 75] (see also Ref. [33] before the DESI data release and earlier studies [79, 80]). Such time-dependent dark energy may be connected to the cosmological constant problem [81, 67], and indeed, it arises naturally in models that address it [81]. In this context, a hilltop axion model with matter effects to also address the $S_{8}$ tension has been proposed [82] (see also [40] for domain wall formation and [83, 84, 85, 86, 87, 88] for general hilltop axion models).
@@ 2 String axion domain walls and isotropic cosmic birefringence
In this section, we investigate a scenario in which an axion initially located near the hilltop of its potential leads to the formation of domain walls without strings, and explore the resulting implications for CB. For our purposes, it is sufficient to consider the following potential:
@@
where $\Lambda$ is a dynamical scale, $m_{\phi}\equiv\Lambda^{2}/f_{\phi}$ is the mass of the axion. The second line represents an expansion around the origin. This potential is invariant under the discrete shift symmetry transformation (1), and consequently admits degenerate vacua. One of the maxima is located at the origin, and there are degenerate vacua at $\phi=\pm\pi f_{\phi},\pm 3\pi f_{\phi},\cdots$. In the following we consider a case in which $\phi$ is initially around the origin so that the dynamical range of $\phi$ is limited to $2\pi f_{\phi}<\phi<2\pi f_{\phi}$.
@@
Note that the onset of oscillations could be delayed due to the flat shape of the potential when $|\phi|\ll f_{\phi}$, namely around the hilltop. This anharmonic effect delays the domain wall formation as discussed below. Suppose that $\phi_{0}$ is sufficiently close to the hilltop at the origin such that
@@
Figure 1: Redshift of domain wall formation estimated using Eq. (11) for $\sigma_{\rm DW}=8m_{\phi}f_{\phi}^{2}=(0.1\,\mathrm{MeV})^{3}$, $(1\,\mathrm{MeV})^{3}$, and $(10\,\mathrm{MeV})^{3}$ from left to right. The horizontal dashed line denotes the recombination epoch, $z\sim 1100$. Given the possible delay due to the anharmonic effect (see Fig. 2), this estimate is regarded as the maximal redshift of the domain wall formation.
Figure 2: Delay of the domain wall formation with hilltop initial conditions for $\sigma_{\rm DW}=(1\,{\rm\,MeV})^{3}$, corresponding to the green line in Fig. 1. Left panel: Evolution of the axion field $\phi$ for $f_{\phi}=10^{15}\,{\rm\,GeV}$ with different initial conditions, $\phi_{\rm ini}/f_{\phi}=10^{-24}$, $10^{-15}$, $10^{-6}$, $10^{-2}$ from bottom to top. These represent typical displacements from the potential maximum, expected to be of order $H_{\rm inf}/f_{\phi}$. The horizontal dashed line denotes $\phi/f_{\phi}=\pi$. Right panel: Redshift at the domain wall formation, $z_{\mathrm{form}}$, defined as the time when $\phi$ first crosses $\pi f_{\phi}$, as a function of $\phi_{\rm ini}/f_{\phi}$. We adopt the decay constants of $f_{\phi}=10^{14}$ GeV (dotted), $3\times 10^{14}$ GeV (dashed), and $10^{15}$ GeV (solid). The solid line corresponds to the left panel.
Domain walls form shortly after the axion begins to oscillate, at the time when condition (11) is satisfied. The corresponding redshift at the onset of oscillations, estimated using this condition, is shown in Fig. 1. Here we use the fact that the domain wall tension is given by $\sigma_{\rm DW}\simeq 8m_{\phi}f_{\phi}^{2}$ for the potential (10) and fix the tension by $\sigma_{\rm DW}=(0.1\,\mathrm{MeV})^{3}$, $(1\,\mathrm{MeV})^{3}$, and $(10\,\mathrm{MeV})^{3}$, from left to right. Once we fix $\sigma_{\mathrm{DW}}$, we can relate $f_{\phi}$ to $m_{\phi}$, and then to the time of the domain wall formation using Eq. (11). Note that this estimate is valid when the field fluctuations are not much smaller than $f_{\phi}$ or when the zero mode is not very close to the hilltop. If the fluctuations are small and the zero mode is in the vicinity of the hilltop, the anharmonic effect becomes significant and delays the onset of oscillations. Fig. 2 illustrates how the anharmonic effect causes this delay. As shown in Fig. 2, the initial field value of the axion must be extremely close to the potential maximum in order to significantly delay the onset of domain wall formation. In this sense, the redshift shown in Fig. 1 should be interpreted as the maximal redshift of the domain wall formation.
@@
where $\bar{\phi}_{\rm LSS}$ denotes the angular average of $\phi_{\rm LSS}(\Omega)$ over the LSS, and we have used the fact that $N_{e}\simeq 50\sim 60$. Since the observer resides in one of the two vacua after domain wall formation, the isotropic CB is given by
@@
where the sign in the second expression is chosen to be positive to match the observed sign of the rotation angle, namely, we take $c_{\gamma}\phi_{\rm today}>0$.
Note that the origin of the isotropic CB in this scenario differs from that in the original KBCB mechanism. There, domain wall formation has already occurred by the time of last scattering, and both adjacent vacua are populated on the LSS. The average axion field is given by $\bar{\phi}_{\rm LSS}=(\phi_{\rm min,1}+\phi_{\rm min,2})/2$, where $\phi_{\rm min,1}$ and $\phi_{\rm min,2}$ denote the field values of the adjacent minima, one of which corresponds to our local vacuum. The isotropic CB then arises from averaging over these domains, while anisotropic CB is also generated due to spatial variations in the field. In this case, if we take into account the population bias between the two vacua, the predicted rotation angle may carry uncertainty depending on the significance of the population bias. In contrast, in the present scenario, the axion remains near the hilltop until well after recombination, and the field on the LSS is nearly homogeneous. Consequently, the CB from CMB photons, which is induced by the difference in $\phi_{\mathrm{LSS}}$ and $\phi_{\mathrm{today}}$, is purely isotropic. Then, the isotropic CB angle is determined by Eq. (15) and is free from ambiguity related to population bias. Remarkably, despite the differences in the underlying dynamics, the predicted isotropic rotation angle is the same in both scenarios.
@@
We also note that in this scenario, the formation time of domain walls can be significantly later than the maximal estimate shown in Fig. 1, if the fluctuations are sufficiently small. This behavior is illustrated in Fig. 2, where we show the evolution of a homogeneous axion field $\phi$ in the left panel and the redshift at the domain wall formation $z_{\mathrm{form}}$ depending on the initial condition $\phi_{\mathrm{ini}}/f_{\phi}$ in the right panel. Here, we numerically solve the equation of motion
@@
The initial field values are taken as $\phi_{\rm ini}=\{10^{-9},\,1,\,10^{9},\,10^{13}\}\,{\rm\,GeV}$, arranged from bottom to top in the plot, with $f_{\phi}=10^{15}\,{\rm\,GeV}$ and $\sigma_{\rm DW}=(1\,{\rm\,MeV})^{3}$. These initial values may be considered to be of order $H_{\rm inf}$. The time of domain wall formation is identified as the moment when the field reaches $\phi/f_{\phi}\sim\pi$, and the oscillations begin.
We see that, in order to avoid the domain wall problem for a string axion, the domain walls must form after the recombination era. As a result, we do not expect anisotropic birefringence in the CMB. However, if the walls form well before reionization, $\phi$ lies in different vacua in different domains at the reionization, and thus photons from that epoch can still experience anisotropic birefringence. As a result, CMB photons from the reionization epoch experience anisotropic CB with a rotation angle comparable to the isotropic CB for photons from the recombination epoch. Since the reionization contribution to the CMB polarization anisotropies is limited, such a signal is more difficult to probe than the anisotropic signal for photons from the recombination epoch [30, 36, 40]. Current observational constraints on such anisotropic CB remain consistent even with scenarios involving cosmic strings while substantial improvement is expected in future full-sky CMB experiments [95].
In addition, if the domain walls formed significantly earlier than the present time, they could have emitted gravitational waves as discussed in Ref. [96]. They also evaluate the spectrum of the gravitational waves and discuss the CMB constraints on such gravitational waves. Such gravitational waves are searched for alongside CMB anisotropies induced by the gravitational potential of domain walls [89, 97, 98, 99].
If the domain walls have only just formed today, they would not have reached the scaling regime. In that case, many closed domain walls could exist within the current horizon. This is possible if the wall thickness is much smaller than the horizon size, which can occur when the onset of oscillations is significantly delayed by the anharmonic effect near the hilltop. Photons from galaxies at relatively low redshift may experience a polarization rotation if they cross domain walls an odd number of times. The resulting rotation angle is predicted to be $0.42\,c_{\gamma}\,\mathrm{deg}$, which is twice the isotropic birefringence angle observed in the CMB.
@@ 3 Time-Varying Dark Energy
If domain walls are formed, however, it becomes challenging to simultaneously satisfy the constraint on their tension—imposed by the requirement that their energy density remain below roughly $10^{-5}$ of the present total energy density—and realize a significant deviation of the dark energy equation of state from that of a cosmological constant (w=1w=-1). Although it is in principle possible to suppress the energy stored in the domain walls while enhancing the energy in the bulk oscillations of the axion field, which we will comment on later, we here focus on the implications of a hilltop initial condition for a string axion—suggested by the isotropic CB—for the DESI DR2 observations, assuming that domain walls are not formed.
@@
[…]
@@
Figure 3: Redshift dependence of the dark energy equation of state parameter $w_{\rm DE}(z)$ in the hilltop axion scenario. The red (blue) line corresponds to $f_{\phi}=10^{16}\,\mathrm{GeV}$, $\phi_{\rm ini}=3\,\mathrm{GeV}$, and $m_{\phi}=1.13\times 10^{-31}\,\mathrm{eV}$ ($f_{\phi}=6\times 10^{16}\,\mathrm{GeV}$, $\phi_{\rm ini}=10^{15}\,\mathrm{GeV}$, and $m_{\phi}=1.89\times 10^{-32}\,\mathrm{eV}$). The black point shows the DESI DR2 result from Refs. [61, 100]. The horizontal dashed line represents $w_{\mathrm{DE}}=-1$, corresponding to the cosmological constant, and the vertical dashed line represents the current time, $z=0$.
@@
where we define $t_{\mathrm{osc}}$ by $t_{\mathrm{osc}}\equiv 2/(3H_{\Lambda\mathrm{CDM}}(z_{\mathrm{osc}}))$ with $H_{\Lambda\mathrm{CDM}}(z_{\mathrm{osc}})\equiv H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z_{\mathrm{osc}})^{3}}$. From these conditions, we obtain the required tuning of the initial condition for fixed values of $f_{\phi}$ (or $m_{\phi}$) as shown in Fig. 4. We also show the parameters from Fig. 3 by red and blue dots in Fig. 4, which roughly match the analytical estimate. See, e.g., Refs. [40, 82, 83, 84, 85, 86, 87, 88] for mechanisms that realize such a hilltop initial condition. For instance, thermal or matter effects could make the potential maximum the minimum temporally in the early universe[40, 82], and the axion can be stabilized there, thereby realizing the hilltop initial condition.